### v1

# Dual Ascent Diffusion for Inverse Problems

Abstract
Ill-posed inverse problems are fundamental in many domains, ranging from astrophysics to medical imaging. Emerging diffusion models provide a powerful prior for solving these problems. Existing maximum-a-posteriori (MAP) or posterior sampling approaches, however, rely on different computational approximations, leading to inaccurate or suboptimal samples. To address this issue, we introduce a new approach to solving MAP problems with diffusion model priors using a dual ascent optimization framework. Our framework achieves better image quality as measured by various metrics for image restoration problems, it is more robust to high levels of measurement noise, it is faster, and it estimates solutions that represent the observations more faithfully than the state of the art.
1 Introduction
We are interested in solving inverse problems, where an unknown image or signal $\mathbf{x}$ is estimated from noisy and corrupted observations $\mathbf{y}$. These types of problems arise across science and engineering, for example, in image restoration [24], astrophysics [1], medical imaging [35], protein structure determination [14, 25, 22], among other domains. In all cases, a linear or nonlinear function $\mathcal{A}\left(\cdot\right)$ models a domain-specific image formation process. Although the likelihood of observations $p\left(\mathbf{y}|\mathbf{x}\right)$ depends on the statistical model of the noise in the observations, closed-form expressions exist for specific cases. For example, the image formation model for zero-mean Gaussian i.i.d. noise with variance $\sigma^{2}$ is $\mathbf{y}=\mathcal{A}\left(\mathbf{x}\right)+\mathcal{N}\left(\mathbf{0},\sigma^{2}\right)$ and its log-likelihood $\log p\left(\mathbf{y}|\mathbf{x}\right)=\frac{-1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}$, up to an additive term that does not depend on $\mathbf{x}$. Given the likelihood of the observations and a prior $p\left(\mathbf{x}\right)$, inverse problem solvers aim at either maximizing or sampling from the posterior $p(\mathbf{x}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{x})p(\mathbf{x})$ in a Bayesian framework. Most inverse problems are ill-posed, making the prior a crucial component of the solution-finding process.
Maximum-a-posteriori (MAP) approaches aim at maximizing the posterior to find the most likely solution given a prior. Early approaches used “hand-crafted” priors to promote smoothness, piece-wise constancy via Total Variation [29, 3], or sparsity in a transform domain [12], while most modern approaches use some form of neural network [40]. The plug-n-play (PnP) approach [37], for example implemented by the Alternating Direction Method of Moments (ADMM) [4] algorithm, is a popular and versatile framework to solve MAP problems by leveraging (Gaussian) denoisers as priors. MAP finds the single, most likely solution to an inverse problem. However, oftentimes one is interested in sampling from the posterior of all feasible solutions. For this reason, many recent works [18, 33, 6, 7, 32, 20, 21, 38, 15, 26, 39] focus on posterior sampling using powerful pretrained diffusion models as priors, as surveyed in [9]. While these recent diffusion posterior sampling methods show great promise, they are all fundamentally limited by the optimization framework that is used to combine the likelihood of the image formation model and the prior during optimization.
In this work, we do not aim to develop a method that provably samples from the posterior, but instead focus on deriving an optimization strategy that accurately and efficiently solves the MAP problem using a prior given by a pretrained diffusion model and a dual ascent–based optimization framework inspired by ADMM [4]. Our approach, dubbed DDiff, is faster and shown to achieve better reconstruction quality compared with the state of the art for image restoration problems, including single-image super resolution, inpainting, deblurring, phase retrieval, and high-dynamic range imaging. Moreover, DDiff is more robust to high levels of measurement noise, and our reconstructions more faithfully model the observations by exhibiting a lower likelihood than existing methods. The latter is important because the log-likelihood of the observations $\textrm{log}\,p\left(\mathbf{y}|\mathbf{x}\right)$ is an indicator for the level of hallucination a generative prior, such as a diffusion model, introduces when computing a solution.
2 Background on Inverse Problems
2.1 Maximum-a-Posteriori Solutions
A maximum-a-posteriori (MAP) solution aims to find the solution $\mathbf{x}_{\textrm{MAP}}$ that maximizes the posterior $p\left(\mathbf{x}|\mathbf{y}\right)\propto p\left(\mathbf{y}|\mathbf{x}\right)p\left(\mathbf{x}\right)$. Typically, this is done by minimizing the negative log-likelihood as
$\mathbf{x}_{\textrm{MAP}}=\operatorname*{argmin}_{\mathbf{x}}-\left(\log p\left(\mathbf{y}|\mathbf{x}\right)+\log p\left(\mathbf{x}\right)\right)=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}-\log p\left(\mathbf{x}\right).$ (1)
The alternating direction method of multipliers (ADMM) [4] is a common approach to solving the MAP problem. ADMM attempts to blend the benefits of dual decomposition and augmented Lagrangian methods for constrained optimization. For this purpose, a slack variable $\mathbf{z}$ is introduced to split the objective function in Eq. 1 into a data fidelity term $-\log p\left(\mathbf{y}|\mathbf{x}\right)$ and the log-prior term $-\log p\left(\mathbf{z}\right)$, subject to $\mathbf{x}=\mathbf{z}$. ADMM then forms the Augmented Lagrangian of the split formulation as
$L_{\rho}\left(\mathbf{x},\mathbf{z},\mathbf{u}\right)=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}-\log p\left(\mathbf{z}\right)+\frac{\rho}{2}\left\|\mathbf{x}-\mathbf{z}+\mathbf{u}\right\|_{2}^{2}-\frac{\rho}{2}\left\|\mathbf{u}\right\|_{2}^{2},$ (2)
where $\mathbf{u}$ is the dual variable and $\rho$ is a hyperparameter that defines the strength of the soft constraints. ADMM then applies an alternating gradient descent approach to minimizing the Augmented Lagrangian, resulting in a set of updates on $\mathbf{x,z,u}$ that are applied in an iterative fashion:
$\mathbf{x}$ $\leftarrow\operatorname*{argmin}_{\mathbf{x}}\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}+\frac{\rho}{2}\|\mathbf{x}-\mathbf{z}+\mathbf{u}\|_{2}^{2}$ (3)
$\mathbf{z}$ $\leftarrow\operatorname*{argmin}_{\mathbf{z}}-\log p\left(\mathbf{z}\right)+\frac{\rho}{2}\|\mathbf{x}-\mathbf{z}+\mathbf{u}\|_{2}^{2}=\mathcal{D}\left(\mathbf{x}+\mathbf{u},\tilde{\sigma}^{2}=\frac{1}{\rho}\right)$ (4)
$\mathbf{u}$ $\leftarrow\mathbf{u}+\mathbf{x}-\mathbf{z}.$ (5)
Here, the $\mathbf{x}$-update is an unconstrained least-squares problem that does not depend on the prior and which often has a closed-form, or at least an efficient, solution. An important insight of plug-and-play image restoration methods [37, 5] is the fact that the $\mathbf{z}$-update (Eq. 4) is a denoising problem on the variable $\mathbf{x}+\mathbf{u}$, which can be solved using any Gaussian denoiser $\mathcal{D}\left(\cdot,\tilde{\sigma}^{2}\right)$ assuming the noise level is $\tilde{\sigma}$.
Dual ascent, for example implemented by ADMM, offers several benefits in traditional optimization, including the ability to leverage convexity in the dual problem, leading to simpler, more efficient, and more robust optimization methods [4].
2.2 Diffusion Models and Diffusion Posterior Sampling
The key insight of diffusion models lies in the fact that one can sample from a target distribution $p_{0}(\mathbf{x})$ by first sampling $\mathbf{x}_{T}$ from another distribution $p_{T}$ that is easy to sample from, e.g., a Gaussian, and iteratively applying a reverse diffusion step of the form
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}+\left(1-{\alpha}_{t}\right)\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)+\sqrt{1-{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ (6)
for $t=T\ldots 1$ to generate an image $\mathbf{x}_{0}$. The reverse diffusion process [2, 16] approximates the inverse trajectories of a corresponding forward diffusion $\mathbf{x}_{t}=\sqrt{\bar{\alpha}}_{t}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$, $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$. Here, we adopt the variance-preserving form of forward and reverse diffusion [17]. The factors $\alpha_{t}$ and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ are derived from the noise schedule of the diffusion model [17]. Importantly, the score network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, defined by parameters $\theta$, is a neural network that approximates the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ [34]. This network is learned from training data in the diffusion model pretraining stage. To sample more efficiently, Denoising Diffusion Implicit Models (DDIM) [31] converts Eq. 6 into a non-Markovian process:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)-\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\,\left(\sqrt{1-\bar{\alpha}_{t}}\cdot\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)\,+\sigma_{t}\epsilon,$ (7)
where $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ and $\sigma_{t}$ is a schedule chosen at inference.
In a posterior sampling problem, we aim at sampling from the posterior $p(\mathbf{x}|\mathbf{y})$. For this purpose, many recent methods follow the approach described above, using the posterior score $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}|\mathbf{y})=\nabla_{\mathbf{x}}\log p(\mathbf{y}|\mathbf{x}_{t}=\mathbf{x})+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ instead of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. The second term is equivalent to the unconditional score of the pretrained diffusion model, but the challenge lies in the first, i.e., the conditional score term. The conditional probability $p(\mathbf{y}|\mathbf{x}_{t})$ can be written as a conditional expectation $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[p(\mathbf{y}|\mathbf{x}_{0})]$, but approximating this expectation with Monte Carlo samples is computationally intractable (see Chung et al. [7], for example). For this reason, existing diffusion posterior sampling methods approximate this conditional distribution with a Dirac delta distribution concentrated on $\mathbf{x}_{t}$ [18] or on $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[\mathbf{x}_{0}]$ [7]. A number of other diffusion posterior sample methods have been proposed [9], each providing a different approximation for the conditional expectation. Most recently, Zhang et al. [39] introduced DAPS (Decoupled Annealing Posterior Sampling), a two-step iterative approach that mitigates the accumulation of errors along the sampling trajectory through an Markov chain Monte Carlo-based equilibration step at $t=0$. DAPS demonstrated state-of-the-art results on both linear and nonlinear inverse problems, but remains limited by the necessity to approximate the conditional probability $p(\mathbf{x}_{0}|\mathbf{x}_{t})$ and by the finiteness of the number of MCMC steps.
3 Method
We derive our approach to solving MAP problems with pretrained diffusion model priors in the following.
3.1 Diffusion Plug-and-Play ADMM for Image Restoration
We dub the naive approach for using ADMM with a pretrained diffusion model Diff-PnP-ADMM. For this purpose, we apply the ADMM framework as discussed in Sec. 2.1 and use the pretrained diffusion model as a one-step denoiser $\mathcal{D}$ in the $\mathbf{z}$-update (i.e., Eq. 4). This is done by applying Tweedie’s formula [13] and replacing the $\mathbf{z}$-update with
$\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}+\mathbf{u}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}+\mathbf{u},t)\right)$ (8)
where $t$ decreases at each step of the ADMM loop, which implicitly constrains the relationship between the diffusion schedule $\alpha_{t}$ and the soft constraint parameter $\rho$.
3.2 Dual Ascent Diffusion (DDiff)
At its core, the ADMM method iterates over 3 steps. The $\mathbf{x}$-update corresponds to a data matching step (Eq. 3), the $\mathbf{z}$-update can be seen as a denoising step (Eq. 4) and the dual update stems from the introduction of the dual variable $\mathbf{u}$ for the constraint $\mathbf{x}=\mathbf{z}$ (Eq. 5). While the $\mathbf{x}$ and dual updates are straightforwardly derived from the original ADMM framework, our methodological contribution primarily consists in showing that a pretrained diffusion model can be used more efficiently in the $\mathbf{z}$-update than the naive approach (Eq. 8).
$\mathbf{x}$-update. In order for the method to be directly applicable to any differentiable forward model $\mathcal{A}$, whether linear or nonlinear, we replace the minimization problem of Eq. 3 with a single gradient step
$\mathbf{x}\leftarrow\mathbf{v}-\gamma\nabla_{\mathbf{v}}\|\mathbf{y}-\mathcal{A}(\mathbf{v})\|_{2}^{2},\quad\mathbf{v}=\mathbf{z}-\mathbf{u},$ (9)
where $\gamma$ is a step size that can be adjusted at each iteration. This approach is also known as linearized ADMM [27].
$\mathbf{u}$-update. The update of the dual variable is readily available in Eq. 5.
$\mathbf{z}$-update. Following Eq. 4, the $\mathbf{z}$-update consists in denoising $\mathbf{x}+\mathbf{u}$ for a certain noise level $\tilde{\sigma}^{2}$. However, it is crucial to note that since the score model $\mathbf{s}_{\theta}(\mathbf{x},t)$ is only trained on points sampled from $p_{t}$, it is a poor approximation of the true score whenever $\mathbf{x}$ is unlikely under $p_{t}$ (i.e., when $p_{t}(\mathbf{x})\ll 1$). In other words, the score model at time $t$ is only accurate on points belonging to the diffusion manifold at time $t$ [8]. Because $\mathbf{x}+\mathbf{u}$ does not in general belong to this manifold, we propose to replace the $\mathbf{z}$-update with
$\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right),$ (10)
where $\mathbf{x}_{t}$ is defined recursively following $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and
$\mathbf{x}_{t-1}\leftarrow\underbrace{\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{x}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\hat{\epsilon}+\sigma_{t}\epsilon}_{\text{DDIM update}}+\underbrace{\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{u}}_{\text{Re-scaled }\mathbf{u}}.$ (11)
In this equation, $\hat{\epsilon}=\left(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\cdot\mathbf{x}\right)/\sqrt{1-\bar{\alpha}_{t}}$ and $\sigma_{t}$ is a hyperparameter. The first part of the right-hand side corresponds to the DDIM update where the “predicted $\mathbf{x}_{0}$” is $\mathbf{x}$. The second part adds the dual variable $\mathbf{u}$, re-scaled to match the signal level of $\mathbf{x}_{t-1}$.
Combined in an iterative fashion, these three steps define our method, described with pseudo-code in Algorithm 1.
Figure 1: Overview of DDiff. This example illustrates the motion deblurring task. Our method alternates between three steps ($\mathbf{x}$-update, $\mathbf{z}$-update, dual update). On the right, we show the evolution of $\mathbf{x}$, $\mathbf{z}$, and $\mathbf{u}$ throughout optimization.
3.3 Comparison to Other Diffusion-Based Variable Splitting Methods
Prior methods, such as DiffPIR [42] and DCDP [23], are MAP-based optimization methods, specifically built upon the half-quadratic splitting (HQS) method, which lacks dual variables in its formulation. Our work naturally extends existing frameworks, transforming them from HQS-style approaches to those using dual variables, including ADMM. By incorporating Lagrange multipliers that accumulate constraint violations across the iterations, DDiff offers improved empirical performance on challenging inverse problems where measurement consistency is crucial. In particular, we note that removing the dual update from DDiff would exactly emulate DiffPIR [42] (with the right choice of $\sigma_{t}$ and $\gamma_{t}$, see supplements).
Algorithm 1 DDiff
0:  $T$, $\mathcal{A}(\cdot)$, $\{\sigma_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, $\mathbf{s}_{\theta}$, $\mathbf{y}$, $\{\gamma_{t}\}_{t=1}^{T}$, $t_{0}$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{u}=\mathbf{0}$.
2:  for $t=T-1$ to 00 do
3:     $\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)$ Denoising step (Eq. 10)
4:     $\mathbf{x}\leftarrow\mathbf{z}-\mathbf{u}-\gamma_{t}\nabla_{\mathbf{v}=\mathbf{z}-\mathbf{u}}\|\mathbf{y}-\mathcal{A}(\mathbf{v})\|^{2}$ Measurement step (Eq. 9)
5:     $\hat{\epsilon}\leftarrow\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\cdot\mathbf{x}\right)$
6:     $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})\>\textbf{if}\>t>t_{0}\>\textbf{else}\>\epsilon=0$
7:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{x}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\hat{\epsilon}+\sigma_{t}\epsilon+\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{u}$ Reverse diffusion (Eq. 11)
8:     $\mathbf{u}\leftarrow\mathbf{u}+\mathbf{x}-\mathbf{z}$Dual update (Eq. 5)
9:  end for
10:  return  $\mathbf{x}_{0}$
4 Experiments
4.1 Experimental Setup
Datasets and metrics.
We evaluate our method on two image datasets, FFHQ $256\times 256$ [19] and ImageNet $256\times 256$ [10]. We utilize pretrained models from [7] on the FFHQ dataset and from [11] on the ImageNet dataset. We randomly selected 100 images from the validation set for both datasets, and the images were normalized to $[-1,1]$. Our main evaluation metrics include peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), learned perceptual image patch similarity (LPIPS) [41], and residual error, defined as $\|\mathbf{y}-\mathcal{A}(\mathbf{x})\|_{2}^{2}-\sigma^{2}$ (giving zero in expectation for an ideal solver).
Inverse problems.
Our method is evaluated on multiple inverse problems. Linear tasks include super-resolution (4×4$\times$4 × downsampling), Gaussian deblurring, motion deblurring, inpainting with a $128\times 128$ box, and inpainting with a random mask that removes $70\%$ of the pixels. Nonlinear tasks consist of phase retrieval (oversampling ratio of $2.0$), for which we report the best result out of five runs due to the intrinsic instability of the task, nonlinear deblurring, and high dynamic range (2×2$\times$2 × dynamic range). All measurements include additive Gaussian noise ($\sigma=0.05$).
Baselines.
We compare against the state of the art: DAPS [39], DPS [7], and DiffPIR [42]. Notably, DiffPIR was not proven to handle nonlinear tasks. We closely evaluate our method against DAPS as it achieves the best results among the baselines.
4.2 Main Results
Quantitative evaluation for the linear and nonlinear tasks on FFHQ and ImageNet datasets are shown in Table 1. Our method outperforms the baselines on the vast majority of the tasks, especially in terms of perceptual similarity and residual error. This is further demonstrated in Fig. 1, where we show a qualitative comparison between the baselines and our method. Overall, DDiff reconstructs finer details with less visual artifacts. More examples of each task and choice of hyperparameters are shown in the supplements.
Moreover, DDiff exhibits significantly increased robustness to higher measurement noise. As shown in Fig. 3, although PSNR may be slightly lower than DAPS in the very low measurement noise region (approximately $\sigma<0.05$), PSNR degrades less rapidly as $\sigma$ increases. LPIPS is generally lower for DDiff than that of DAPS in all levels of noise. LPIPS follows a similar trend, especially on the phase retrieval task, where DDiff divides LPIPS by 3 in the high-noise region ($\sigma=0.3$). To visualize the reconstruction quality of DDiff compared to DAPS at the extreme case of measurement noise level ($\sigma=0.3$), refer to Fig. 4. These results suggest that DDiff could be more appropriate for solving challenging inverse problems where measurement noise is prominent, such as but not limited to low-dose CT reconstruction, cryo-electron microscopy, and sparse-view tomography.
Table 1: Quantitative evaluation. Comparing different methods for 5 linear and 3 nonlinear tasks on FFHQ and ImageNet datasets. This evaluation uses 100 validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. All tasks are run with a noise of standard deviation $\sigma=0.05$.
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× DDiff (ours) 30.07 0.824 0.211 0.0028 25.81 0.656 0.396 0.0038 DAPS 29.34 0.783 0.190 0.0029 25.44 0.636 0.295 0.2871 DPS 24.42 0.486 0.346 0.0050 21.10 0.351 0.408 0.0052 DiffPIR 23.71 0.440 0.423 0.0087 20.75 0.312 0.517 0.0093 Inpainting (Box) DDiff (ours) 24.88 0.831 0.110 0.0077 21.15 0.743 0.240 0.0119 DAPS 24.52 0.742 0.174 0.0099 21.22 0.714 0.230 0.2050 DPS 24.68 0.810 0.079 0.0033 19.63 0.725 0.254 0.0412 DiffPIR 19.02 0.527 0.252 0.0106 16.02 0.520 0.329 0.0132 Inpainting (Random) DDiff (ours) 33.08 0.877 0.050 0.0205 28.39 0.758 0.136 0.0241 DAPS 30.76 0.801 0.156 0.0293 27.32 0.725 0.189 0.0788 DPS 30.79 0.807 0.083 0.0217 27.31 0.737 0.235 0.0980 DiffPIR 18.53 0.362 0.622 0.0264 15.82 0.191 0.842 0.1080 Gaussian Deblurring DDiff (ours) 28.87 0.800 0.119 0.0026 22.29 0.471 0.415 0.0046 DAPS 29.63 0.789 0.177 0.0027 25.90 0.658 0.269 0.2373 DPS 27.77 0.704 0.140 0.0029 23.07 0.628 0.332 0.0084 DiffPIR 26.16 0.624 0.297 0.0031 21.64 0.393 0.497 0.0093 Motion Deblurring DDiff (ours) 28.24 0.785 0.129 0.0058 24.16 0.585 0.242 0.0079 DAPS 29.17 0.797 0.186 0.0059 26.61 0.710 0.241 0.1825 DPS 27.93 0.714 0.130 0.0061 23.36 0.611 0.321 0.0082 DiffPIR 22.01 0.327 0.499 0.0074 18.93 0.248 0.586 0.0084 Phase Retrieval DDiff (ours) 29.94 0.816 0.120 1.4690 18.54 0.494 0.262 1.4659 DAPS 29.60 0.768 0.182 1.4770 20.23 0.449 0.397 1.4685 DPS 22.24 0.540 0.307 1.5412 16.03 0.396 0.444 1.5540 DiffPIR 10.04 0.036 0.783 1.6711 9.61 0.021 0.794 1.5410 Nonlinear Deblurring DDiff (ours) 31.48 0.873 0.120 0.0027 29.68 0.805 0.207 0.0035 DAPS 28.45 0.764 0.188 0.0042 27.28 0.718 0.213 0.2643 DPS 25.39 0.643 0.258 0.0095 28.42 0.691 0.271 0.0053 DiffPIR 19.79 0.331 0.583 0.0273 22.13 0.459 0.435 0.0167 High Dynamic Range DDiff (ours) 26.05 0.873 0.129 0.0459 26.50 0.800 0.108 0.0541 DAPS 27.39 0.846 0.163 0.0505 26.10 0.825 0.171 0.7172 DPS 25.79 0.793 0.165 0.0734 22.72 0.721 0.273 0.4951 DiffPIR 17.69 0.645 0.296 0.1292 18.23 0.637 0.289 0.1105
Figure 2: Qualitative results. DDiff demonstrates sharper and cleaner results compared to DPS and DAPS. All tasks are run with a noise of standard deviation $\sigma=0.05$.
(a) Gaussian deblurring
(b) Phase retrieval
Figure 3: Effect of measurement noise level. DDiff demonstrates greater robustness as noise increases. This evaluation uses 10 FFHQ validation images.
Figure 4: Qualitative results at high measurement noise level. We use 10 randomly-chosen validation images and show the results of phase retrieval at $\sigma=0.3$, comparing our method to the state-of-the-art algorithm DAPS [39].
4.3 Ablation Studies
Evaluation of time efficiency and quality of samples.
One of the most limiting factors of sampling speed for diffusion-based methods is the number of neural function evaluations (NFEs). It measures the number of times the underlying neural network is called during computation, which in our case is the number of times we call the score model during inference. Therefore, we assess the average sampling time and sample quality of our method and of existing baselines as a function of the number of NFEs in Fig. 5. For the same number of NFEs, DDiff achieves better perceptual quality and faster sampling speed compared to DAPS. The enhanced computational efficiency is a consequence of two key factors: reduced backpropagation requirements for gradient calculations and the absence of supplementary MCMC procedures within sampling iterations, both of which are present in the DAPS framework.
(a) Gaussian deblurring
(b) Phase retrieval
Figure 5: Evaluation of time efficiency and quality of samples. The y-axis shows LPIPS value and the x-axis shows the time (in sec.) taken to generate one sample image on a GeForce RTX 2080 Ti 12GB GPU. The number after the method name (500, 2k, etc.) indicates the NFEs. This evaluation uses 100 FFHQ validation images.
Effect of dual variable and noising step.
To illustrate the importance of dual variable $\mathbf{u}$ and the additional noising step (Eq. 11) in our algorithm, we conduct an experiment to compare DDiff (with noise, with $\mathbf{u}$) to three other variants: Diff-PnP-HQS (no noise, no $\mathbf{u}$), Diff-PnP-ADMM (no noise, with $\mathbf{u}$), and DDiff-HQS (with noise, no $\mathbf{u}$) in Table 2. Our analysis indicates that the dual variable implementation alone decreases performance, as evidenced by Diff-PnP-ADMM’s inferior metrics compared to Diff-PnP-HQS. Without the noising step, the dual variable introduces high-frequency artifacts that compromise the diffusion model’s efficacy. However, when incorporated alongside the noising step, the dual variable yields a substantial enhancement in performance, as evidenced by the superior quantitative metrics achieved by DDiff compared to DDiff-HQS. Our experimental results confirm that both components—dual variable and noising step—must be implemented together, as demonstrated by the significant performance improvement observed when transitioning from Diff-PnP-HQS to DDiff.
Table 2: Ablating the dual variable and noising step. We compare DDiff (with noise, with $\mathbf{u}$) to other variants, including Diff-PnP-HQS (no noise, no $\mathbf{u}$), Diff-PnP-ADMM (no noise, with $\mathbf{u}$), and DDiff-HQS (with noise, no $\mathbf{u}$). All methods are compared in a fair manner by using the same DDIM noise schedule. This evaluation uses 10 FFHQ validation images and reports the average metric value. The best and second-best results are distinguished by bold and underlined marks, respectively. DDiff, with the dual variable and noising step, significantly outperforms other methods.
Super Resolution 4× Inpainting (Box) Inpainting (Random) Gaussian Deblurring Method PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ DDiff (ours) 30.10 0.821 0.199 0.0034 26.01 0.855 0.087 0.0032 33.21 0.883 0.043 0.0028 29.08 0.801 0.112 0.0025 DDiff-HQS 25.45 0.558 0.369 0.0057 19.01 0.566 0.221 0.0091 15.37 0.254 0.790 0.0167 26.32 0.642 0.267 0.0031 Diff-PnP-ADMM 13.79 0.434 0.572 0.0304 12.72 0.546 0.486 0.0298 16.79 0.540 0.402 0.0087 6.18 0.077 0.820 0.2552 Diff-PnP-HQS 14.04 0.447 0.535 0.0287 12.61 0.557 0.476 0.0271 15.85 0.513 0.431 0.0168 5.82 0.036 0.858 0.2812 Motion Deblurring Phase Retrieval Nonlinear Deblurring High Dynamic Range Method PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. PSNR SSIM LPIPS Res. $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ DDiff (ours) 28.14 0.779 0.112 0.0026 30.58 0.834 0.096 0.0028 31.25 0.867 0.111 0.0033 26.95 0.877 0.109 0.0072 DDiff-HQS 21.75 0.396 0.397 0.0036 12.86 0.114 0.679 0.0057 13.51 0.186 0.543 0.1252 17.76 0.672 0.284 0.1260 Diff-PnP-ADMM 6.35 0.089 0.810 0.1840 12.84 0.266 0.637 0.0081 5.23 0.007 0.830 0.2463 7.85 0.126 0.875 0.4797 Diff-PnP-HQS 6.39 0.091 0.813 0.1752 13.63 0.287 0.597 0.0077 6.02 0.086 0.827 0.2537 7.56 0.098 0.885 0.5287
5 Discussion
In this work, we introduce Dual ascent for solving Diffusion model-based inverse problems (DDiff), a method achieving better quantitative and qualitative reconstruction results than the state of the art, especially at high measurement noise levels. Our method combines optimization in primal and dual spaces and appropriately modifies the $\mathbf{z}$-update of the ADMM loop to efficiently leverage a pretrained diffusion model as a prior, achieving improved perceptual quality while improving measurement consistency and reducing runtime, making DDiff especially suitable for solving large-scale, complex inverse problems.
Limitations and Future Work.
While our work demonstrates promising results for pixel-space diffusion models, an extension to latent diffusion models would be an interesting avenue of future work [28, 30]. Furthermore, extending our approach to higher-dimensional data, such as 3D or video representations presents valuable directions with additional challenges.
Conclusion.
With our work, we contribute to the growing community efforts on solving inverse problems robustly, with strong adherence to the observations and using state-of-the-art diffusion-based priors.
Broader Impact Statement
The study conducted in this paper introduces a novel approach to leveraging generative models for solving inverse problems. Potential applications of our work include both image restoration and image editing. We do not condone using our work with the intent to generate harmful or deceiving contents, or spread misinformation.
Appendix A Technical Appendices and Supplementary Material
A.1 Statistical Significance
Building on the main results table, we report 95% confidence intervals for DDiff and DAPS in Table S1, highlighting the statistical significance of our improvement, compared to DAPS [39].
Table S1: Quantitative evaluation with confidence intervals. We show the average metric values over 100 validation images and the corresponding 95% confidence intervals for DDiff and DAPS.
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× DDiff (ours) 30.07$\pm$0.41 0.824$\pm$0.008 0.211$\pm$0.009 (2.85±$\pm 0.06$)$\cdot 10^{-3}$ 25.81$\pm$0.72 0.656$\pm$0.029 0.396$\pm$0.029 (3.83±$\pm 0.29$)$\cdot 10^{-3}$ DAPS 29.34$\pm$0.33 0.783$\pm$0.006 0.190$\pm$0.006 (2.87±$\pm 0.06$)$\cdot 10^{-3}$ 25.44$\pm$0.56 0.636$\pm$0.017 0.295$\pm$0.011 (2.87±$\pm 0.36$)$\cdot 10^{-1}$ Inpainting (Box) DDiff (ours) 24.88$\pm$0.50 0.831$\pm$0.005 0.110$\pm$0.005 (7.75±$\pm 9.21$)$\cdot 10^{-3}$ 21.15$\pm$0.68 0.743$\pm$0.009 0.240$\pm$0.011 (1.19±$\pm 0.29$)$\cdot 10^{-2}$ DAPS 24.52$\pm$0.40 0.742$\pm$0.006 0.174$\pm$0.006 (9.86±$\pm 6.76$)$\cdot 10^{-3}$ 21.22$\pm$0.70 0.714$\pm$0.007 0.230$\pm$0.011 (2.05±$\pm 0.24$)$\cdot 10^{-1}$ Inpainting (Random) DDiff (ours) 33.08$\pm$0.37 0.877$\pm$0.004 0.050$\pm$0.003 (2.05±$\pm 0.15$)$\cdot 10^{-2}$ 28.39$\pm$0.76 0.758$\pm$0.019 0.136$\pm$0.018 (2.41±$\pm 0.35$)$\cdot 10^{-2}$ DAPS 30.76$\pm$0.27 0.801$\pm$0.005 0.156$\pm$0.003 (2.93±$\pm 0.16$)$\cdot 10^{-2}$ 27.32$\pm$0.63 0.725$\pm$0.013 0.189$\pm$0.010 (7.88±$\pm 0.70$)$\cdot 10^{-2}$ Gaussian Deblurring DDiff (ours) 28.87$\pm$0.43 0.800$\pm$0.010 0.119$\pm$0.005 (2.60±$\pm 0.07$)$\cdot 10^{-3}$ 22.29$\pm$0.84 0.471$\pm$0.039 0.415$\pm$0.035 (4.68±$\pm 0.39$)$\cdot 10^{-3}$ DAPS 29.63$\pm$0.36 0.789$\pm$0.006 0.177$\pm$0.005 (2.67±$\pm 0.06$)$\cdot 10^{-3}$ 25.90$\pm$0.64 0.658$\pm$0.020 0.269$\pm$0.011 (2.37±$\pm 0.31$)$\cdot 10^{-1}$ Motion Deblurring DDiff (ours) 28.24$\pm$0.38 0.785$\pm$0.009 0.129$\pm$0.005 (5.82±$\pm 0.27$)$\cdot 10^{-3}$ 24.16$\pm$0.65 0.585$\pm$0.027 0.242$\pm$0.014 (7.97±$\pm 0.60$)$\cdot 10^{-3}$ DAPS 29.17$\pm$0.36 0.797$\pm$0.007 0.186$\pm$0.005 (5.91±$\pm 0.28$)$\cdot 10^{-3}$ 26.61$\pm$0.65 0.710$\pm$0.020 0.241$\pm$0.011 (1.82±$\pm 0.25$)$\cdot 10^{-1}$ Phase Retrieval DDiff (ours) 29.94$\pm$0.88 0.816$\pm$0.025 0.120$\pm$0.019 1.469$\pm$0.002 18.54$\pm$1.23 0.494$\pm$0.045 0.262$\pm$0.029 1.465$\pm$0.002 DAPS 29.60$\pm$0.88 0.768$\pm$0.020 0.182$\pm$0.018 1.477$\pm$0.002 20.23$\pm$1.27 0.449$\pm$0.044 0.397$\pm$0.032 1.469$\pm$0.002 Nonlinear Deblurring DDiff (ours) 31.48$\pm$0.29 0.873$\pm$0.005 0.120$\pm$0.006 (2.74±$\pm 0.07$)$\cdot 10^{-3}$ 29.68$\pm$0.62 0.805$\pm$0.016 0.207$\pm$0.019 (3.55±$\pm 0.18$)$\cdot 10^{-3}$ DAPS 28.45$\pm$0.38 0.764$\pm$0.007 0.188$\pm$0.006 (4.16±$\pm 0.43$)$\cdot 10^{-3}$ 27.28$\pm$0.62 0.718$\pm$0.017 0.213$\pm$0.011 (2.64±$\pm 0.33$)$\cdot 10^{-1}$ HDR DDiff (ours) 26.05$\pm$0.68 0.873$\pm$0.007 0.129$\pm$0.009 (4.59$\pm$0.27)$\cdot 10^{-2}$ 26.50$\pm$0.56 0.800$\pm$0.018 0.108$\pm$0.014 (5.41$\pm$0.27)$\cdot 10^{-2}$ DAPS 27.39$\pm$0.56 0.846$\pm$0.010 0.163$\pm$0.012 (5.05±$\pm 0.18$)$\cdot 10^{-2}$ 26.10$\pm$0.79 0.825$\pm$0.021 0.171$\pm$0.019 (7.17±$\pm 0.75$)$\cdot 10^{-1}$
A.2 Relating DDiff and DiffPIR
Here we show the equivalence of DDiff and DiffPIR (see Algorithm S1) under specific parameter choices. DiffPIR introduces $\zeta$ as a hyperparameter controlling the stochasticity of the algorithm and $\bar{\sigma}_{t}$ as a parameter controlling the strength of the likelihood step. Our algorithm is equivalent to the linearized version of DiffPIR when:
  1. i.
    $\sigma_{t}=\sqrt{\zeta\cdot(1-\bar{\alpha}_{t-1})}$;
  2. ii.
    $\gamma_{t}=\frac{\bar{\sigma}_{t}^{2}}{2\lambda\sigma^{2}}$ where $\sigma$ is the measurement noise standard deviation;
  3. iii.
    We remove early the relaxation parameter $t_{0}$;
  4. iv.
    We remove the dual updates.
Algorithm S1 DiffPIR
0:  $T$, $\mathcal{A}(\cdot)$, $\{\bar{\sigma}_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, $\mathbf{s}_{\theta}$, $\mathbf{y}$, $\zeta$, $\lambda$, $\sigma$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
2:  for $t=T-1$ to 00 do
3:     $\mathbf{x}_{0}^{(t)}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)$
4:     $\hat{\mathbf{x}}_{0}^{(t)}\leftarrow\mathbf{x}_{0}^{(t)}-\frac{\bar{\sigma}_{t}^{2}}{2\lambda\sigma^{2}}\nabla_{\mathbf{x}_{0}^{(t)}}\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0}^{(t)})\|^{2}$
5:     $\hat{\epsilon}\leftarrow\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{x}}_{0}^{(t)}\right)$
6:     $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
7:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{\mathbf{x}}_{0}^{(t)}+\sqrt{1-\bar{\alpha}_{t-1}}(\sqrt{1-\zeta}\hat{\epsilon}+\sqrt{\zeta}\epsilon_{t})$
8:  end for
9:  return  $\mathbf{x}_{0}$
A.3 DDiff Variants Algorithms
This section presents the precise algorithms corresponding to the three variants of DDiff discussed in the ablation study. Note that DDiff-HQS S4 is equivalent to DiffPIR S1 if we satisfy (i.-iii.) from section A.2.
Algorithm S2 Diff-PnP-HQS (no noise, no $\mathbf{u}$)
0:  $T$, $\mathcal{A}(\cdot)$, $\{\sigma_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, $\mathbf{s}_{\theta}$, $\mathbf{y}$, $\{\gamma_{t}\}_{t=1}^{T}$, $t_{0}$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
2:  for $t=T-1$ to 00 do
3:     $\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)$
4:     $\mathbf{x}_{t-1}\leftarrow\mathbf{z}-\gamma_{t}\nabla_{\mathbf{z}}\|\mathbf{y}-\mathcal{A}(\mathbf{z})\|^{2}$
5:  end for
6:  return  $\mathbf{x}_{0}$
Algorithm S3 Diff-PnP-ADMM (no noise, with $\mathbf{u}$)
0:  $T$, $\mathcal{A}(\cdot)$, $\{\sigma_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, $\mathbf{s}_{\theta}$, $\mathbf{y}$, $\{\gamma_{t}\}_{t=1}^{T}$, $t_{0}$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{u}=\mathbf{0}$.
2:  for $t=T-1$ to 00 do
3:     $\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+\mathbf{u}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t}+\mathbf{u},t)\right)$
4:     $\mathbf{x}_{t-1}\leftarrow\mathbf{z}-\mathbf{u}-\gamma_{t}\nabla_{\mathbf{v}=\mathbf{z}-\mathbf{u}}\|\mathbf{y}-\mathcal{A}(\mathbf{v})\|^{2}$
5:     $\mathbf{u}\leftarrow\mathbf{u}+\mathbf{x}_{t-1}-\mathbf{z}$
6:  end for
7:  return  $\mathbf{x}_{0}$
Algorithm S4 DDiff-HQS (with noise, no $\mathbf{u}$)
0:  $T$, $\mathcal{A}(\cdot)$, $\{\sigma_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, $\mathbf{s}_{\theta}$, $\mathbf{y}$, $\{\gamma_{t}\}_{t=1}^{T}$, $t_{0}$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
2:  for $t=T-1$ to 00 do
3:     $\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)$
4:     $\mathbf{x}\leftarrow\mathbf{z}-\gamma_{t}\nabla_{\mathbf{z}}\|\mathbf{y}-\mathcal{A}(\mathbf{z})\|^{2}$
5:     $\hat{\epsilon}\leftarrow\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\cdot\mathbf{x}\right)$
6:     $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})\>\textbf{if}\>t>t_{0}\>\textbf{else}\>\epsilon=0$
7:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{x}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\hat{\epsilon}+\sigma_{t}\epsilon$
8:  end for
9:  return  $\mathbf{x}_{0}$
A.4 Inverse Problems Setup
Most inverse problems are implemented using the same approach described in [7], except for the HDR task which follows the setup in [39]. We set a fixed random seed for inpainting, motion deblurring, and nonlinear deblurring for fair comparison. Specific parameters are defined as follows:
  • Super resolution: 4×4$\times$4 × downsampling factor
  • Inpainting: $128\times 128$ box mask and $70\%$ random mask
  • Gaussian & motion deblurring: $61\times 61$ kernel size with standard deviations of $3.0$ and $0.5$, respectively
  • Phase retrieval: oversampling with ratio $k/n$ where $k=2$ and $n=8$
  • Nonlinear deblurring: blur kernel generated using [36]
  • HDR: 2×2$\times$2 × dynamic range
A.5 Hyperparameter Choices
The step size $\gamma_{t}$ for the measurement step is defined by a step function below:
$\gamma_{t}=\gamma_{0}\cdot f(t_{\gamma})$ where $f(t_{\gamma})=\begin{cases}3.3&\text{for }t>t_{\gamma},\\ 0.1&\text{for }t\leq t_{\gamma}.\end{cases}$
To enhance sample quality, the parameter $\gamma_{t}$ is reduced at time step $t_{\gamma}$. This adjustment encourages the traversal to relax toward the prior manifold while diverging from the likelihood manifold. As a result, the generated samples exhibit reduced noise and improved visual fidelity.
We used two different choices of $\sigma_{t}$ in the main experiments.
  • $\sigma_{t}=\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$ (this choice allows the generative process to become a DDPM) was used for Gaussian deblurring, motion deblurring, nonlinear deblurring, and HDR.
  • $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t-1}}$ was used for phase retrieval, inpainting (box and random), and super resolution.
We report the hyperparameters used in the main experiments in Table S2 and S3.
Table S2: DDiff hyperparameter settings for different tasks on FFHQ.
Super Res. 4×4$\times$4 × Inpaint (Box) Inpaint (Rand.) Gaussian deblur
$\gamma_{0}$ 18 30 50 2.9
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 1 1 1 50
Motion deblur Phase retrieval Nonlinear deblur HDR
$\gamma_{0}$ 2.9 38 2.5 3.5
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 80 1 120 120
Table S3: DDiff hyperparameter settings for different tasks on ImageNet.
Super Res. 4×4$\times$4 × Inpaint (Box) Inpaint (Rand.) Gaussian deblur
$\gamma_{0}$ 18 50 50 1.8
$t_{\gamma}$ 90 500 90 90
$t_{0}$ 1 1 1 50
Motion deblur Phase retrieval Nonlinear deblur HDR
$\gamma_{0}$ 1.5 38 2.5 3.8
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 80 1 120 100
A.6 More Qualitative Results
We provide additional qualitative results in Fig. S1 to S8.
Figure S1: Visual comparison of DDiff and baselines on super resolution task.
Figure S2: Visual comparison of DDiff and baselines on box inpainting task.
Figure S3: Visual comparison of DDiff and baselines on random inpainting task.
Figure S4: Visual comparison of DDiff and baselines on Gaussian deblurring task.
Figure S5: Visual comparison of DDiff and baselines on motion deblurring task.
Figure S6: Visual comparison of DDiff and baselines on phase retrieval task.
Figure S7: Visual comparison of DDiff and baselines on nonlinear deblurring task.
Figure S8: Visual comparison of DDiff and baselines on HDR task.